A water tank can be filled by a gasoline fueled water pump in 28 minutes and by an electric pump in 21 minutes. How long would it take to fill the tank if both pumps were used?
Answer by Cromlix(4381)

Gas fuelled water pump takes 28 minutes
Electric pump takes 21 minutes
Gas pump takes 1/28 job/min
Electric pump takes 1/21 job/min
Together rate = 1/x job/min
1/28 + 1/21 = 1/x
Common denominator = 588x
21x + 28x = 588
49x = 588
x = 12
So together both pumps will
fill the tank in 12 minutes.
Hope this helps.
:-)